OpenStudy (anonymous):

Area of a shaded region, posted below. I have some, dont know where to proceed.

OpenStudy (anonymous):

 \[\int\limits_{1/2}^{2} 1/x \]

OpenStudy (anonymous):

thats all i have at the moment

OpenStudy (anonymous):

the anti derivative of 1/x = ln(x) so you evaluate this from 1/2 to 2 so its ln(2) - ln(1/2)

OpenStudy (anonymous):

thank you so much, i missed my anti-derivative class so i ont know when to use it

OpenStudy (anonymous):

the answer is 1.39

OpenStudy (anonymous):

np glad to help

OpenStudy (anonymous):

Your instructor might prefer it as \(\ln 4\), which is exact.